# Temperature-Dependent Calibration Procedures for the Silicon Photomultiplier Readout of the Cosmic Ray Veto Detector for the Mu2e Experiment

### v1

@@ 6.1.3 SiPM Light Yield Dependence on Over-Voltage
The SiPM response to the flasher LED light versus bias voltage, for runs taken under the same temperature, is shown in Fig. 15, left. It shows a linear behavior. The SiPM response at a fixed bias voltage of 54.0 V versus temperature is shown in Fig. 15. That dependence is also linear, the average change in light yield being 0.08)%/K. Convoluting the temperature dependence of the breakdown voltage with the light yield dependence on the breakdown voltage gives the light yield dependence with respect to changes in the breakdown voltage: (21.82.1)%/V. This is parameter in Table 1.
There are limitations to this method. Due to different normalization, the calculations are specific to the choice of bias voltage, which has been somewhat arbitrarily selected. At a fixed bias voltage, the SiPM’s light yield dependence on the temperature is a manifestation of the over-voltage dependence of SiPM’s photon detection efficiency and cross-talk. Due to the SiPM-to-SiPM difference in the breakdown voltage, the over-voltage and hence the normalization are not exactly uniform across all the SiPMs. In addition, the light emission of the LED used as the flasher also had a temperature dependence. Although the flasher was placed outside the temperature chamber, the external LED temperature varied slightly with the chamber temperature. This introduced a systematic uncertainty in the above analysis. The effect is relatively small, as Sec. 7 will prove. Hence, the corrections using the parameters derived above work well. Readers wishing to make a more precise light yield correction can choose to produce a light yield versus over-voltage curve for a specific SiPM and apply the correction accordingly.
@@ 6.2.1 Bias Supply FEB Temperature Dependence
This correction accounts for temperature variations on the FEB board that affect the biases supplied to SiPMs. This effect arises from the temperature sensitivity of the resistors used to provide the bias to the SiPMs.131313In the upcoming FEB version, we plan to incorporate resistors with reduced temperature coefficients. We measured the temperature coefficient of the FEB provided biases by placing the board in the environmental chamber and measuring the voltages at the SiPMs using a hand-held multi-meter at various chamber temperatures. On average the bias voltage increases by as the FEB temperature increases (Fig. 17). This is parameter in Table 1.
@@ 6.2.2 AFE Chip Gain FEB Temperature Dependence
Figure 18 shows the SPE values of four example SiPMs at different FEB temperatures, using the setup illustrated in Fig. 16. Notre that there were minor SiPM temperature differences, which were corrected as discussed in Sec. 6.1. On average, a slope of was observed. This change is from the temperature dependence of the SiPM bias provided by the FEB, as well as the temperature dependence of the AFE amplifier of the FEB. To determine the effect of the bias change, and remove it, we take determined above and multiply it by parameter of Table 1, 3 ADCns/V, to get . Subtracting this off gives : the temperature dependence of the AFE amplifier, which is the only other gain-related temperature dependence. This is parameter from Table 1. Note that this value is consistent with the information from the AFE data sheet. However, it should also be noted this value is specific to the gain setting we use for the AFE chip; a different setting will lead to a different correction.
@@ 6.3 Parameter Validation
To verify the calibration technique described above, in this section we first perform a sanity check using the same data from which the parameters were derived. In the following section we apply the technique to cosmic-ray data taken with the CRV modules at Wideband.
We reconstruct the SPE before and after the corrections, interpolating the SPE at a nominal bias of 54.0 V (corresponding to an over-voltage of 2.5 V at C) for different temperature runs. The before and after results are shown in Fig. 19 for the first four SiPMs. Before the temperature corrections, SPE temperature dependence was , averaging over four SiPMs chosen at random. After corrections, the temperature dependence was reduced to .
We also measured the LED light yield at 54.0 V bias voltage after the corrections, as shown in Fig. 20. This can be compared with the right plot in Fig. 15, which shows the response of the same SiPMs before any corrections. The correction works well, and the LED light yields across a 50∘C temperature range show no temperature dependence difference at a level below 5%. Note that there are small SiPM-to-SiPM differences, and the deviations from perfectly flat lines in Fig. 20 come from nonlinearity in the relation between SiPM light yield and over-voltage, as well as variations in the LED light emission: the determination of the parameter assumes the light from the LED was constant.141414In our measurements, the LED temperature was controlled to better than a 10∘C range, corresponding to a light emission difference of less than 0.5%. To further improve this measurement, one should keep the LED and the SiPMs monitoring its light emission at a constant temperature. Nonetheless, these deviations are comparably small. As shown in the next section, the correction scheme we derived here is sufficient for the typical data we collect since the temperature variations at the storage facility, and the expected variations during Mu2e operations, are much smaller.151515Temperature measurements in the detector hall, without the top shielding installed and with doors to the detector hall occasionally open to the outside, have a maximum yearly variation of 6∘C. That variation is expected to be much less when the detector hall is closed for data taking.
@@ 7 Validation of Calibration Scheme with CRV Modules
Modules at the Wideband detector hall are stored in stacks as shown in Fig. 21. Three plastic scintillator trigger paddles read out with photomultiplier tubes, were placed at one meter from the readout end, perpendicular to the direction of the CRV counters. A 3-fold coincidence of signals in the trigger paddles within 100 ns produced a trigger; primarily cosmic ray muons. The trigger was distributed to the FEBs through the ROC.
@@
The two runs taken with comparable ambient conditions, characterized by SiPM temperatures of approximately C, resulted in light yield differences of %. The third run, performed at approximately C, was compared to one of the C runs. The uncorrected and temperature-corrected photoelectron yields are shown in Fig. 22. With a temperature difference of C between the two runs, our correction algorithms yield a light yield difference of % when averaged over all 64 SiPMs.
As a final validation of this method, we show the light yield response to cosmic-ray muons for an example CRV SiPM over three years in Fig. 23. Without temperature corrections, the light yield varies inversely with the temperature. After the correction, the variations are less and there is a clear decline in the light yield with time, or aging of the scintillator. A complete analysis of CRV aging results will be presented in a forthcoming presentation.
Currently, the temperature at the Mu2e hall has a maximum yearly variation of C. When buttoned up for data taking it is expected to have less variation during Mu2e operations with the shielding in place and the doors to the detector hall closed. The stable operation environment combined with the temperature correction scheme described above will ensure a consistent performance of the CRV system.
@@ 8 Summary
A method for calibration and handling the temperature dependence of non-temperature controlled SiPMs used for the Mu2e cosmic ray veto detector (CRV) has been described in detail. The method corrects the single photoelectron measurement to a reference temperature to allow comparison of measurements taken at different temperatures. For this to work the temperature dependence of all of the components affecting the readout of the SiPMs must be well understood: hence, the temperature dependence measurements were made in an environmental chamber. The method was validated using an independent experimental setup with cosmic-ray muon data collected using CRV modules. When the results from data samples were taken from two runs with large temperature differences (C) the corrected results were found to be consistent, that is no residual dependence on temperature was observed.

### v2

@@ 6.1.3 SiPM Light Yield Dependence on Over-Voltage
The SiPM response to the flasher LED light versus bias voltage, for runs taken under the same temperature, is shown in Fig. 15 left, resulting in a linear behavior. The SiPM response at a fixed bias voltage of 54.0 V versus temperature is shown in Fig. 15 right. That dependence is also linear, the average change in light yield being 0.08)%/K. Convoluting the temperature dependence of the breakdown voltage with the light yield dependence on the breakdown voltage gives the light yield dependence with respect to changes in the breakdown voltage: (21.82.1)%/V. This is parameter in Table 1.
There are limitations to this method. Due to different normalization, the calculations are specific to the choice of bias voltage, which has been somewhat arbitrarily selected. At a fixed bias voltage, the SiPM’s light yield dependence on the temperature is a manifestation of the over-voltage dependence of SiPM’s photon detection efficiency and cross-talk. Due to the SiPM-to-SiPM difference in the breakdown voltage, the over-voltage and hence the normalization are not exactly uniform across all the SiPMs. In addition, the light emission of the LED used as the flasher also had a temperature dependence. Although the flasher was placed outside the temperature chamber, the external LED temperature varied slightly with the chamber temperature. This introduced a systematic uncertainty in the above analysis. The effect is relatively small, as proven in Sec. 7. Hence, the corrections using the parameters derived above work well. Readers wishing to make a more precise light yield correction can choose to produce a light yield versus over-voltage curve for a specific SiPM and apply the correction accordingly.
@@ 6.2.1 Bias Supply FEB Temperature Dependence
This correction accounts for temperature variations on the FEB board that affect the biases supplied to SiPMs. This effect arises from the temperature sensitivity of the resistors used to provide the bias to the SiPMs.151515In the upcoming FEB version, we plan to incorporate resistors with reduced temperature coefficients. We measured the temperature coefficient of the FEB provided biases by placing the board in the environmental chamber and measuring at various chamber temperatures the voltages at the SiPMs using a hand-held multi-meter. On average the bias voltage increases by as the FEB temperature increases (Fig. 17). This is parameter in Table 1.
@@ 6.2.2 AFE Chip Gain FEB Temperature Dependence
Figure 18 shows the SPE values of the four example SiPMs at different FEB temperatures, using the setup illustrated in Fig. 16. Note that there were minor SiPM temperature differences, which were corrected as discussed in Sec. 6.1. On average, a slope of was observed. This change is from the temperature dependence of the SiPM bias provided by the FEB, as well as the temperature dependence of the AFE amplifier of the FEB. To determine the effect of the bias change, and remove it, we take determined above and multiply it by parameter of Table 1, 3 ADCns/V, to get . Subtracting this off gives : the temperature dependence of the AFE amplifier, which is the only other gain-related temperature dependence. This is parameter from Table 1. Note that this value is consistent with the information from the AFE data sheet. However, it should also be noted this value is specific to the gain setting we use for the AFE chip; a different setting will lead to a different correction.
@@ 6.3 Parameter Validation
To verify the calibration technique described above, in this section we first perform a validation check using the same data from which the parameters were derived. In the following section we apply the technique to cosmic-ray data taken with the CRV modules at Wideband.
We reconstruct the SPE before and after the corrections, interpolating the SPE at a nominal bias of 54.0 V (corresponding to an over-voltage of 2.5 V at C) for different temperature runs. The before and after results are shown in Fig. 19 for the four SiPMs. Before the temperature corrections, SPE temperature dependence was , averaging over the four arbitrarily chosen SiPMs. After corrections, the temperature dependence was reduced to .
We also measured the LED light yield at 54.0 V bias voltage after the corrections, as shown in Fig. 20. This can be compared with the right plot in Fig. 15, which shows the response of the same SiPMs before any corrections. When averaged over all channels, the residual slope is (0.10.1)%, verifying that any remaining temperature dependence is small. The correction works well, and the LED light yields across a 50∘C temperature range vary by less than 5%. Note that there are small SiPM-to-SiPM differences, and the deviations from perfectly flat lines in Fig. 20 come from nonlinearity in the relation between SiPM light yield and over-voltage, as well as variations in the LED light emission: the determination of the parameter assumes the light from the LED was constant.161616In our measurements, the LED temperature was controlled to better than a 10∘C range, corresponding to a light emission difference of less than 0.5%. To further improve this measurement, one should keep the LED and the SiPMs monitoring its light emission at a constant temperature. Nonetheless, these deviations are comparably small. As shown in the next section, the correction scheme we derived here is sufficient for the typical data we collect since the temperature variations at the storage facility, and the expected variations during Mu2e operations, are much smaller.171717Temperature measurements in the detector hall, without the top shielding installed and with doors to the detector hall occasionally open to the outside, have a maximum yearly variation of 6∘C. That variation is expected to be much less when the detector hall is closed for data taking.
@@ 7 Validation of Calibration Scheme with CRV Modules
Modules at the Wideband detector hall are stored in stacks as shown in Fig. 21. Three plastic scintillator trigger paddles read out with photomultiplier tubes, were placed at one meter from the readout end, perpendicular to the length of the CRV counters. A 3-fold coincidence of signals in the trigger paddles within 100 ns produced a trigger; primarily cosmic ray muons. The trigger was distributed to the FEBs through the ROC.
@@
The two runs taken with comparable ambient conditions, characterized by SiPM temperatures of approximately C, resulted in light yield differences of %. The third run, performed at approximately C, was compared to one of the C runs. The uncorrected and temperature-corrected photoelectron yields are shown in Fig. 22. With a temperature difference of C between the two runs, our correction algorithms resulted in a light yield difference of % when averaged over all 64 SiPMs.
As a final validation of this method, we show the light yield response to cosmic-ray muons for an example CRV SiPM over three years in Fig. 23. Without temperature corrections, the light yield varies inversely with the temperature. After the correction, the variations are mitigated and there is a clear decline in the light yield with time, or aging of the scintillator. A complete analysis of CRV aging results will be presented in a forthcoming presentation.
Currently, the temperature at the Mu2e hall has a maximum yearly variation of C. When the detector hall is closed up for data taking it is expected to have less variation during Mu2e operations with the shielding in place and the doors to the detector hall closed. The stable operation environment combined with the temperature correction scheme described above will ensure a consistent performance of the CRV system.
@@ 8 Summary
A method for calibration and handling the temperature dependence of non-temperature controlled SiPMs used for the Mu2e cosmic ray veto detector (CRV) has been described in detail. The method corrects the single photoelectron measurement to a reference temperature to allow comparison of measurements taken at different temperatures. For this to work, the temperature dependence of all of the components affecting the readout of the SiPMs must be well understood: hence, the temperature dependence measurements were made in an environmental chamber. The method was validated using an independent experimental setup with cosmic-ray muon data collected using CRV modules. When the results from data samples were taken from two runs with large temperature differences (C) the corrected results were found to be consistent, that is no residual dependence on temperature was observed.